OpenStudy (anonymous):

Find the distance betweent the pairs (6,8) and (18,17)

OpenStudy (anonymous):

\[D=\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\]

Directrix (directrix):

15 http://easycalculation.com/analytical/distance.php

OpenStudy (anonymous):

distance between (6,8) and (18,17): (17 - 8)² + ( 18 -6)² = 81+ 144 = √ 225 = 15